OpenStudy (anonymous):

Find f '(x) and f ''(x). f(x)= x/9-e^x

OpenStudy (anonymous):

really having an issue with differentiables containing square roots and e^x

OpenStudy (anonymous):

the derivative of e^x is always e^x, forever, it's its own derivative

OpenStudy (anonymous):

what if a number is attached like the 9-e^x?

OpenStudy (anonymous):

derivativave of 9 is 0,

OpenStudy (anonymous):

ahh so you do seperate it... the derivative of 9 then e^x

OpenStudy (shadowfiend):

So, when you take a derivative, you always take the separate derivative of parts that are added or subtracted together. For example: f(x) = g(x) + h(x) f'(x) = g'(x) + h'(x) f''(x) = g''(x) + h''(x) And: f(x) = g(x) - h(x) f'(x) = g'(x) - h'(x) f''(x) = g''(x) - h''(x)

OpenStudy (anonymous):

right

OpenStudy (shadowfiend):

So let's take your problem in particular: \[f(x) = \frac{x}{9} - e^x\] You have to first take the derivative of \(\frac{x}{9}\), then that of \(e^x\). Then, you can subtract the latter from the former.

OpenStudy (anonymous):

\[(x(e ^{x}))-(1(9+e^{x}))/(9+e^{x})^{2}\]

OpenStudy (shadowfiend):

Is that a different problem?

OpenStudy (anonymous):

its not subtract its divide

OpenStudy (anonymous):

x divided by 9+e^{x}

OpenStudy (shadowfiend):

Ah, got it.

OpenStudy (shadowfiend):

So: \[f(x) = \frac{x}{9 + e^x}\]?

OpenStudy (anonymous):

yes and the other was the answer i got

OpenStudy (shadowfiend):

I've always found it easiest to use product rule when I have division like that. In this case, you're looking at: \[f(x) = x(x + e^x)^{-1}\]

OpenStudy (shadowfiend):

So in this case, we have product rule. Product rule says: f(x) = h(x)g(x) f'(x) = h'(x)g(x) + h(x)g'(x)

OpenStudy (anonymous):

and thats the answer to fprime?

OpenStudy (anonymous):

sorry to be the pain! lol im a non traditional student that never took calc in high school sad thing is im trying to major in it lmao

OpenStudy (anonymous):

what is wrong with the answer i found?

OpenStudy (shadowfiend):

Sorry, browser retardedness.

OpenStudy (anonymous):

i feel ya

OpenStudy (shadowfiend):

Also, what I meant was that you're looking at : \[f(x) = x(9 + e^x)^{-1}\]

OpenStudy (shadowfiend):

Looks like the only issue you had was a sign issue.

OpenStudy (shadowfiend):

I ultimately got: \[\frac{9 + e^x - xe^x}{(9 + e^x)^2}\]